# FREE: The Foundational Semantic Recognition for Modeling Environmental Ecosystems

### v4

@@ Abstract
Modeling environmental ecosystems is critical for the sustainability of our planet, but is extremely challenging due to the complex underlying processes driven by interactions amongst a large number of physical variables. As many variables are difficult to measure at large scales, existing works often utilize a combination of observable features and locally available measurements or modeled values as input to build models for a specific study region and time period. This raises a fundamental question in advancing the modeling of environmental ecosystems: how to build a general framework for modeling the complex relationships amongst various environmental data over space and time? In this paper, we introduce a framework, FREE, which maps available environmental data into a text space and then converts the traditional predictive modeling task in environmental science to a semantic recognition problem. The proposed framework leverages recent advances in Large Language Models (LLMs) to supplement the original input features with natural language descriptions. This framework facilitates capturing the data semantics and allows harnessing the irregularities of input features. When used for long-term prediction, FREE has the flexibility to incorporate newly collected observations to enhance future prediction. The efficacy of FREE is evaluated in the context of two societally important real-world applications, predicting stream water temperature in the Delaware River Basin and predicting annual corn yield in Illinois and Iowa. Beyond the superior predictive performance over multiple baselines, FREE is shown to be more data- and computation-efficient as it can be pre-trained on simulated data generated by physics-based models.
@@ 1 Introduction
Understanding the dynamics of environmental ecosystems is critical for the sustainable management of natural resources and mitigating natural disasters such as algal blooms and floods. This is becoming even more important given the compelling need for food and water supply from a growing world population and a more unpredictable climate Mbow et al. (2020); Konar et al. (2011). Modeling environmental ecosystems is challenging as these systems are shaped by the complex interactions of a large number of physical variables, such as weather, soil, water, and plants. Hence, it often requires the combination of data from the network of weather stations, remote sensing, and field measurements to jointly model complex system dynamics. However, many physical variables cannot be easily measured over a large scale. They are often only available at certain locations or during certain time periods, or they are derived from models based on established physical theories. Even for those observable variables, they are often sparse due to the substantial cost required for the data collection (e.g., sensor deployment and field survey).
@@
This paper focuses on two important applications, predicting stream water temperature and annual crop yield. Success in both of these prediction tasks can assist a range of applications in optimizing resource allocation and management strategies. For example, the changing climate is making it harder to grow crops in the US Corn Belt, threatening food supplies and farmer livelihoods; drinking water reservoir operators in the Delaware River Basin (our study region) need to supply safe drinking water to over 15 million people while also maintaining sufficient cool water temperatures in the river network downstream of the reservoirs Williamson et al. (2015). Given the importance of these two tasks in ensuring food and water security, scientists from multiple domains, including hydrology, agronomy, meteorology, and environmental engineering, have been developing physics-based models (also referred to as process-based models) to simulate underlying physical processes for different components of aquatic and agricultural ecosystems. Despite their extensive use, the majority of physics-based models are necessarily approximations of reality due to incomplete knowledge or excessive complexity in modeling certain processes. More recently, machine learning (ML) models have been increasingly considered as an alternative approach for modeling environmental ecosystems given their computational efficiency in the prediction phase and their ability to automatically extract complex data patterns. ML methods have found success in many environmental applications, including climate science Rolnick et al. (2022), hydrology Chen et al. (2022), and agriculture Liu et al. (2022).
@@
However, current ML approaches face several major challenges in fully leveraging available data for modeling ecosystems. First, ML models often treat input physical variables as independent numerical features but do not explicitly capture the nature of these variables and their interdependent physical and ecological relationships. This is further exacerbated by the scarcity of observation data in many real-world environmental ecosystems, which limits the ability of ML models to automatically extract generalizable relationships between features. Second, traditional ML approaches are unable to harness inconsistent feature inputs. Current studies on local environmental ecosystems often enhance the prediction of their target study region by leveraging multiple types of input features related to the target variable. However, the selected features often vary between different studies. Concretely, besides meteorological data (e.g., solar radiation) that are commonly used to drive physics-based models, prior works also explored including other measurements on the environment and target system (e.g., soil properties Chen et al. (2023), land use and geometric structures Cho and Kim (2022)), proxy observations derived from remote sensing Mkhabela et al. (2011), and other physical variables simulated by physics-based models Jia et al. (2021b). These features may not always be available for every data sample, which poses a challenge in training a global model that can utilize different input features. Finally, there is a lack of a general pipeline for dynamically incorporating auxiliary observations into the current methods, hindering their ability to achieve better performance in long-term prediction. Current works on assimilating auxiliary observations rely on task-specific learning mechanisms or model structures, e.g., Kalman filtering for incorporating new observations Zwart et al. (2023), and graph convolution and invertible network layers for assimilating observations from neighboring samples Chen et al. (2021a). Yet these approaches can be computationally expensive to implement when we have a large state space and/or non-linear system dynamics.
@@
In this paper, we propose a novel method, Foundational semantic Recognition for modeling Environmental Ecosystems (FREE) to address these limitations. The key idea of FREE is to translate the heterogeneous input data into natural language descriptions using large language models (LLMs) and then estimate the target variable through semantic recognition in the text space. This enables addressing the challenges in utilizing different data sources by only manipulating the text space while maintaining the same predictive modeling component (i.e., semantic recognition). In particular, the translation process uses the available features of each data point with allowing different or incomplete input features for different data points. Moreover, it can easily incorporate auxiliary observations (e.g., newly collected observations from the previous day) into the textual description with a properly designed prompt. For semantic recognition, we utilize a separate language model to embed the obtained textual description and then create additional network layers (e.g., long-short term memory (LSTM)) to capture data dependencies. The use of the language model on textual descriptions enables a better capture of the nature and semantics of input features. To further enhance the embedding performance of the language model on environmental descriptions, we pre-train the semantic recognition component using abundant simulated samples generated by physics-based models. This pre-training process also helps the model better learn the general physical relationships encoded in the physics-based models and mitigate the challenge posed by the sparse observations.
@@
We evaluate the proposed method on two real-world datasets: (1) stream water temperature data from the Delaware River Basin, and (2) corn yield data in Illinois and Iowa. Both of them cover a diverse set of locations and long time periods. The results demonstrate the superior predictive performance of the proposed FREE method over multiple baselines, especially with sparse observations. The method is shown to be able to effectively handle different input features and incorporate auxiliary observations. The pre-training process also helps improve the prediction performance of the model when adapted to different locations.
2 Related Work
Physics-based models have been widely used to simulate complex environmental processes, including different components in water cycles and plant growth in agroecosystems Zhou et al. (2021); Srinivasan et al. (2010). However, these models are necessarily approximations of reality and often rely on approximations and parameterizations Gupta and Nearing (2014). Recent advancements in ML offer new opportunities for modeling environmental ecosystems using the rapidly growing earth observation data and ground observations about water, plants, soils, and climate. In particular, prior research has shown the promise of ML-based approaches in modeling agroecosystems Jia et al. (2019); Liu et al. (2022); Ghazaryan et al. (2020) and freshwater ecosystems Rahmani et al. (2021).
Data assimilation has been widely used for assimilating auxiliary observations to optimally adjust the model state of a system. There are many data assimilation approaches used for modeling environmental ecosystems, and they differ in their assumptions about the distributions of the model, data predictions, and speed of computation. The Kalman filter, which iteratively estimates the next state through a forward process and then updates the state using new observations, is a popular data assimilation technique. Many variations of the Kalman filter have also been implemented Evensen and others (2009). Despite their extensive use, these approaches can be computationally expensive when we have a large state space and/or non-linear system dynamics. Recently, researchers have started to use neural networks as an alternative way for implementing data assimilation Brajard et al. (2020); Chen et al. (2021b). Note that this is different from online incremental learning in that model parameters remain the same but only states are changed. The intuition of data assimilation is to adjust the model state when they are disturbed by external factors that are not captured by input features.
Figure 1: Workflow of FREE. Input features are first transformed into natural language descriptions. These descriptions are then processed by a separate language model (LM) to generate embeddings, which are fed to an LSTM layer for making predictions. Simulated data generated by a physics-based model are used to pre-train the LM and LSTM layers, followed by fine-tuning with true observations of the target variable.
To mitigate the need for large training data in ML models, one promising direction is to transfer knowledge embodied in physical simulations to ML models. Simulated data can be generated by physics-based models that represent underlying physical processes using a series of mathematical equations. For example, prior work has shown that ML models pre-trained using simulated data can perform much better under data-scarce scenarios for modeling freshwater and agricultural ecosystems Jia et al. (2021a); Read et al. (2019).
Recent advancements in LLMs have demonstrated remarkable performance in solving language tasks based on human instructions Touvron et al. (2023); Chiang et al. (2023) and have been applied to geoscience applications Mai et al. (2023). More recently, these models have been applied to other types of data. For example, LLMs have been used to textualize tabular information and handle question-answering tasks Borisov et al. (2022); Zhao et al. (2023). Textualized tabular data offers the advantage of being able to handle inputs with different feature sets and is more robust in dealing with missing values.
@@ 3 Method
In this section, we define the problem and introduce our proposed framework. The objective is to predict the target variable $y$ (e.g., water temperature or crop yield) at multiple locations $i\in\{1,...,N\}$ and over a period of time. For clarity, we denote by $\mathbf{x}_{i,t}$ the input features at the location $i$ on date $t$. The input features contain the set of meteorological variables (e.g., solar radiation, rainfall) that are commonly used as drivers for physics-based models. Besides, we include some other physical variables estimated by physics-based models (e.g., cloud cover fraction, groundwater properties) or obtained from other data sources (e.g., soil properties). Notably, these features may be absent for certain locations or time steps. The observations of the target variable at each location $i$ over multiple time steps is $y_{i}=\{{y}_{i,1},{y}_{i,2},...,{y}_{i,T}\}$ which are only sparsely available for certain time steps and locations.
As outlined in Fig. 1, the proposed framework FREE converts the traditional predictive modeling task in environmental science to the semantic recognition problem in a text space created by LLMs (e.g., GPT). A clear benefit of this framework is its ability to harness inputs of different feature sets and incorporate auxiliary information, e.g., new observations collected from the previous day. We also introduce additional model components (e.g., BERT and LSTM) to embed the text data and capture the temporal dependencies. To mitigate the need for large training samples for different ecosystems, we propose to pre-train the model using abundant simulated data generated by existing physics-based models. In the following, we will describe these components in detail.
Figure 2: Input data conversion. Green arrows indicate FREE handling inputs of diverse feature sets with three major inputs in the prompt: prefix $p$, linearized data, and suffix $s$. Red arrow suggests that traditional ML models need separate preprocessing methodologies (e.g., imputation) to address data irregularities effectively.
@@ 3.1 Overall architecture
The proposed FREE framework (shown in Fig. 1) consists of two major components in its architecture.
Input data conversion: To address inconsistencies in the feature set and incorporate auxiliary observations, we propose to transform the original data sample into a corresponding natural language description. This approach facilitates the handling of diverse and potentially incomplete feature sets for different data points, enabling a uniform textual representation of data across varying input scenarios.
Specifically, we leverage existing LLM GPT-3.5 to transform each data point, represented as $\mathbf{x}_{i,t}$, into clear, natural language descriptions $d_{i,t}$. To effectively communicate with GPT-3.5, we use a linearization technique Wang et al. (2023) to construct prompts that consist of a context-setting prefix, the linearized data input, and a directive suffix. The prefix ($p$) provides the model with a background of the dataset, while the suffix ($s$) instructs the model on how to format its output. The complete prompt is thus formulated as:
@@
As depicted in Fig. 2, we format the input features $\mathbf{x}_{i,t}$ of $K$ dimensions into a sequence pairing column names ${c}^{k}_{i,t}$ with the corresponding feature values $\mathbf{x}^{k}_{i,t}$ (e.g., { [rainfall: 0], [solar radiation: 151.14154] } ). This process can be expressed as follows:
@@
Given the prompt, the LLM can generate feature summaries that are both descriptive and succinct, such as: ”On December 4, 2006, there was no recorded rainfall in the Delaware River Basin. The average air temperature was -3.364 degrees Celsius, where freezing-thawing and phase change may occur. The solar radiation measured was 108.26 watts per square meter.”.
The obtained textual descriptions, alongside the observations in the original dataset, form paired data samples as $\{d_{i,t},y_{i,t}\}$, which are then used for tuning the semantic recognition component (to be discussed later).
In summary, the proposed input conversion process helps generate an understandable and flexible representation of input features. More importantly, the use of LLM enables supplementing the raw feature values with semantically meaningful text descriptions and interpreting underlying physical phenomena, which facilitates subsequent modeling components to capture the complex nature and the interactions of input features.
Semantic recognition: Once obtaining the textual descriptions $d_{i,t}$, we utilize a separate language model (LM) $f$ to process the obtained text data and embed them as
@@
In this work, we use the DistilBERT model Sanh et al. (2019) in this step. It is vital to note that the embeddings $\mathbf{e}_{i,t}$ are generated independently across different data points. To effectively capture the intricate data dependencies, we introduce additional network layers that combine multiple data embeddings. In particular, this work uses LSTM layers to capture common temporal data dependencies in the environmental ecosystem, e.g., seasonal changes, and the effect of heavy rainfall on soil water in the next few days. We will explore other advanced network structures, e.g., graph neural networks for modeling streams Jia et al. (2021b) and crop yield Fan et al. (2022), as future work.
@@ 3.2 Handling diverse inputs and auxiliary data
One major goal of modeling environmental ecosystems is long-term prediction in large regions. There are opportunities to enhance ML models towards this goal from two aspects. First, predictive models often have larger predictive errors as time progresses in the test phase due to the accumulated bias in the model state. One could explore leveraging newly collected observations during the testing phase (e.g., the observations collected at time $t$) to mitigate the current model bias and improve the future prediction (e.g., the prediction at $t+1$). Second, for some data samples, we may have access to additional features such as local soil properties or modeled variables from physics-based models. These additional features can be combined with the original features to improve the model prediction. Under the proposed FREE framework, we can make use of the additional observations and features by simply modifying the prompt while keeping the predictive model (i.e., the semantic recognition component) the same.
3.2.1 Incorporating auxiliary observations
In this study, we consider incorporating two types of auxiliary observations to enhance the prediction of the data sample $\mathbf{x}_{i,t}$: (1) the newly collected observation from the previous time $t-1$ and at the same location $i$, and (2) the newly collected observations from the neighboring locations of the target location $i$. We incorporate the information from auxiliary data by updating the linearized data input to include it.
As observations are collected from different dates (e.g., the previous day) with the existing features in $\mathbf{x}_{i,t}$, we need to explicitly include the exact date information for the auxiliary observations and original features. If we consider the auxiliary observations being collected at the date $t-1$ from both the current location $i$ and its neighbor $j\in\mathcal{N}(i)$, the updated linearization process can be expressed as
@@ 3.2.1 Incorporating auxiliary observations
$=$ $[\text{date}:t-1]\cup[{c}_{i}^{y}:y_{i,t-1}]\cup[{c}^{y}_{j}:y_{j,t-1}]$
$\cup[\text{date}:t]\cup\{[{c}^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]\}_{k=1}^{K},$
[…]
It is noteworthy that the auxiliary observations may not always be available. When creating the linearized data, we include the columns of the auxiliary observations only if they are available and skip them otherwise.
3.2.2 Handling different input features
FREE can also handle different input features by using only the available features to create the linearized data while skipping missing features. In particular, if a feature $k$ is not available for the data sample $\mathbf{x}_{i,t}$, then the pair $[c^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]$ will be skipped and not included in the input to the LLM. This entails the LLM generating natural descriptions that focus only on the available features.
The proposed approach enables the use of a combination of heterogeneous data sources for modeling complex environmental ecosystems. It also allows the subsequent semantic recognition to proceed seamlessly on the generated text without the need for manual adjustments to account for data irregularities.

### v5

@@ Abstract
Modeling environmental ecosystems is critical for the sustainability of our planet, but is extremely challenging due to the complex underlying processes driven by interactions amongst a large number of physical variables. As many variables are difficult to measure at large scales, existing works often utilize a combination of observable features and locally available measurements or modeled values as input to build models for a specific study region and time period. This raises a fundamental question in advancing the modeling of environmental ecosystems: how to build a general framework for modeling the complex relationships among diverse environmental variables over space and time? In this paper, we introduce a framework, FREE, that enables the use of varying features and available information to train a universal model. The core idea is to map available environmental data into a text space and then convert the traditional predictive modeling task in environmental science to a semantic recognition problem. Our evaluation on two societally important real-world applications, stream water temperature prediction and crop yield prediction, demonstrates the superiority of FREE over multiple baselines, even in data-sparse scenarios.
@@ 1 Introduction
Understanding the dynamics of environmental ecosystems is critical for the sustainable management of natural resources and mitigating natural disasters such as algal blooms and floods, especially given the compelling need for food and water supply from a growing world population and a more unpredictable climate. Modeling environmental ecosystems is challenging as these systems are shaped by the complex interactions of a large number of physical variables, such as weather, soil, water, and plants. Hence, it often requires the combination of data from the network of weather stations, remote sensing, and field measurements to jointly model complex system dynamics. However, many physical variables are often only sparsely available at certain locations or during certain time periods due to the substantial cost required for the data collection.
@@
We focus on two important applications: predicting stream water temperature and annual crop yield under changing weather conditions. Success in both prediction tasks can support a range of applications in optimizing resource allocation and management strategies. For example, water reservoir operators in the Delaware River Basin (DRB) need to maintain water temperature within a desired range to ensure safe drinking water for over 15 million people while also preserving sufficiently cool water for aquatic life in downstream areas (Williamson et al., 2015); the changing climate is making it making it increasingly difficult to sustain high crop yields in the U.S. Corn Belt, threatening food supplies and farmer livelihoods. While physics-based models have been developed to simulate underlying physical processes for these ecosystems (Zhou et al., 2021), the majority of them are necessarily approximations of reality due to incomplete knowledge or excessive complexity in modeling certain processes (Gupta & Nearing, 2014). Machine learning (ML) models offer an alternative, given their computational efficiency and ability to automatically extract complex data patterns (He et al., 2023; Rolnick et al., 2022; Jia et al., 2019).
@@
However, traditional ML approaches face several major challenges in fully leveraging available data for modeling ecosystems: (1) treating physical variables as independent numerical features without explicitly capturing their interdependent physical and ecological relationships. This is further exacerbated by the scarcity of observation data in many real-world ecosystems, which limits the ability of ML models to automatically extract generalizable relationships between features. (2) inability to harness inconsistent feature inputs. Current studies on local environmental ecosystems often enhance the prediction of their target study region by leveraging multiple types of input features related to the target variable, and the selected features often vary between different studies. Besides meteorological data (e.g., solar radiation) that are commonly used, prior works also explored including other measurements on the environment system (e.g., soil properties (Chen et al., 2023), land use and geometric structures (Cho & Kim, 2022)), and other physical variables simulated by physics-based models (Jia et al., 2021). These features may not always be available for data samples collected from different locations and time periods, which poses a challenge in training a global model that can utilize different input features. (3) lack of a general pipeline that can dynamically incorporate auxiliary observations into the ML model. Current works on assimilating auxiliary observations rely on task-specific learning mechanisms or model structures, e.g., Kalman filtering for incorporating new observations (Zwart et al., 2023), and graph convolution and invertible network layers for assimilating observations from neighboring samples (Chen et al., 2021; Brajard et al., 2020). Yet these approaches can be computationally expensive to implement when we have a large state space and/or non-linear system dynamics. Recent LLM advances show promise for addressing these limitations (Mai et al., 2023). Originally developed for language tasks, LLMs now handle tabular data by converting it to text (Wang et al., 2023; Zhao et al., 2023), enabling flexible feature handling and robustness to missing values.
@@
In this paper, we propose a novel method, Foundational semantic Recognition for modeling Environmental Ecosystems (FREE) to address these limitations. The key idea of FREE is to translate the heterogeneous input data into natural language descriptions using large language models and then estimate the target variable through semantic recognition in the text space. This addresses the challenges in utilizing different data sources by only manipulating the text space while maintaining the same predictive modeling component (i.e., semantic recognition). In particular, the translation process uses the available features for each data point, allowing the set of features to vary across data points. Moreover, the translation process can easily incorporate auxiliary observations (e.g., newly collected observations from the previous day) into the textual description with a properly designed prompt. After obtaining the textual description, we build the semantic recognition component by leveraging a separate language model to embed the text data and incorporating additional network layers (e.g., long-short term memory (LSTM)) to capture temporal dependencies. The use of the language model on textual descriptions enables a better understanding of the nature and semantics of input features. To further enhance the embedding performance of the language model (LM) on environmental descriptions, we pre-train the semantic recognition component using abundant simulated samples generated by physics-based models. This pre-training process also helps the model learn the general physical relationships encoded in the physics-based models and mitigate the challenge posed by the sparse observations.
@@
We evaluate FREE on two real-world tasks: daily stream water temperature prediction in the DRB, and annual county-wise corn yield prediction in Illinois and Iowa. Both of them cover a diverse set of locations and long time periods. FREE shows its superior predictive performance over multiple baselines, especially with sparse observations, and it can effectively handle different input features and incorporate auxiliary observations. The pre-training process also helps improve the prediction performance of the model when adapted to different locations. This work also opens new opportunities to leverage LLMs as a universal tool for general environmental modeling tasks, which are currently often addressed in a fragmented and localized manner due to variations in data collection and processing pipelines.
@@ 2 Method
Figure 1: The framework of FREE: Input features are first transformed into natural language descriptions by LLMs. These descriptions are then processed by a LM to generate embeddings, which are fed to an LSTM layer for making predictions. Simulated labels generated by a physics-based model are used to pre-train the LM and LSTM layers, followed by fine-tuning with true observations of the target variable for enhanced predictions.
In this section, we define the problem and introduce our proposed framework. The objective is to predict the target variable $y$ (e.g., water temperature or crop yield) at multiple locations $i\in\{1,...,N\}$ and over a period of time. For clarity, we denote by $\mathbf{x}_{i,t}$ the input features at the location $i$ on date $t$. The input features contain the set of meteorological variables (e.g., solar radiation, rainfall) that are commonly used as drivers for physics-based models. Besides, we include some other physical variables estimated by physics-based models (e.g., cloud cover fraction, groundwater properties) or obtained from other data sources (e.g., soil properties). Notably, these features may be absent for certain locations or time steps. We represent the observations of the target variable at each location $i$ over multiple time steps as $y_{i}=\{{y}_{i,1},{y}_{i,2},...,{y}_{i,T}\}$. The observations can be only sparsely available for certain time steps and locations.
FREE converts the traditional predictive modeling task in environmental science to the semantic recognition problem in a text space created by LLMs (e.g., GPT). A clear benefit of this framework is its ability to harness inputs of different feature sets and incorporate auxiliary information. We also introduce additional model components to embed the text data and capture the temporal dependencies. To mitigate the need for large training samples for different ecosystems, we pre-train the model using abundant simulated data generated by physics-based models. In the following, we will describe these components in detail.
@@ 2.1 Overall architecture
The FREE framework (shown in Fig. 1) consists of two major components in its architecture.
Input data conversion: To address inconsistencies in the feature set and incorporate auxiliary observations, we propose to transform the original data sample into a corresponding natural language description. This approach facilitates the handling of diverse and potentially incomplete feature sets for different data points, enabling a uniform textual representation of data across varying input scenarios. Specifically, we leverage an existing LLM to transform each data point $\mathbf{x}_{i,t}$ into clear, natural language descriptions $d_{i,t}$. To effectively communicate with the LLM, we use a linearization technique (Wang et al., 2023) to construct prompts that consist of a context-setting prefix, the linearized data input, and a directive suffix. The prefix ($p$) provides the model with a background of the dataset, while the suffix ($s$) instructs the model on how to format its output. The complete prompt is formulated as:
@@
Figure 2: Green arrows indicate FREE handling inputs of diverse feature sets that use linearized data. Red arrow suggests that traditional ML models might need separate preprocessing methodologies to address data irregularities effectively.
As depicted in Fig. 2, we format the input features $\mathbf{x}_{i,t}$ of $K$ dimensions/variables into a sequence pairing column names ${c}^{k}_{i,t}$ with the corresponding feature values $\mathbf{x}^{k}_{i,t}$ (e.g., { [rainfall: 0], [solar radiation: 151.14] } ). This process can be expressed as follows:
@@
Given the prompt, LLM can generate feature summaries that are both descriptive and succinct. , such as: ”On December 4, 2006, there was no recorded rainfall in the Delaware River Basin. The average air temperature was -3.36 degrees Celsius, where freezing-thawing and phase change may occur. The solar radiation measured was 108.26 watts per square meter.”. The obtained textual descriptions, alongside the observations in the original dataset, form paired data samples as $\{d_{i,t},y_{i,t}\}$, which are then used for tuning the semantic recognition component (to be discussed later).
In summary, the proposed input translation process helps generate an understandable and flexible representation of input features. More importantly, the use of LLM enables supplementing the raw feature values with semantically meaningful text descriptions and interpreting underlying physical phenomena, which facilitates subsequent modeling components to capture the complex nature and the interactions of input features.
Semantic recognition: Once obtaining the textual descriptions $d_{i,t}$, we utilize a separate LM $f$ to process the obtained text data and embed them as
@@
In this work, we use the DistilBERT model (Sanh et al., 2019) in this step. It is vital to note that the embeddings $\mathbf{e}_{i,t}$ are generated independently across different data points. To effectively capture the intricate data temporal dependencies, we introduce additional network layers that combine multiple data embeddings. In particular, this work uses LSTM layers to capture common temporal data dependencies in the environmental ecosystem, e.g., seasonal changes, and the effect of heavy rainfall on soil water in the next few days.
@@ 2.2 Handling diverse inputs and auxiliary data
One major goal of modeling environmental ecosystems is long-term prediction in large regions. FREE provides opportunities to enhance ML models towards this goal from two aspects. First, predictive models often have larger predictive errors as time progresses in the test phase due to the accumulated bias in the model state. One could explore leveraging newly collected observations during the testing phase (e.g., the observations collected at previous time t1t-1) to mitigate the current model bias and improve the future predictions (e.g., the prediction after $t$). Second, when training the model across space, some areas may offer more features, such as certain variables related to local soil properties while other regions may only provide weather-related features. Such discrepancies in feature availability pose challenges for building a universal model across space. Under the FREE framework, we can leverage additional observations and handle different features through simple modification of the prompt while keeping the predictive model (i.e., the semantic recognition component) unchanged.
Incorporating auxiliary observations: In this study, we consider incorporating two types of auxiliary observations to enhance the prediction of the data sample $\mathbf{x}_{i,t}$: the newly collected observation from the previous time t1t-1 and at the same location $i$, and from the neighboring locations of the target location $i$. We incorporate the auxiliary data by updating the linearized data input to include it. As observations are collected from different dates with the existing features in $\mathbf{x}_{i,t}$, we need to explicitly include the exact date information for the auxiliary observations and original features. If we consider the auxiliary observations being collected at the date t1t-1 from both the current location $i$ and its neighbor $j\in\mathcal{N}(i)$, the updated linearization process can be expressed as:
@@
$=$ $[\text{date}:t-1]\cup[{c}_{i}^{y}:y_{i,t-1}]\cup[{c}^{y}_{j}:y_{j,t-1}]\cup[\text{date}:t]\cup\{[{c}^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]\}_{k=1}^{K},$
where $c_{i}^{y}$ denotes the column name for the observed labels, $c_{j}^{y}$ denotes the column name for the observed labels from the neighboring location $j$, and $\cup$ indicates concatenation operation across the sequences. LLM then follows the instructions by $p$ and $s$ to generate the description using these data. It is noteworthy that the auxiliary observations may not always be available. When creating the linearized data, we include the columns of the auxiliary observations only if they are available, and skip them otherwise.
Handling different input features: FREE handles different input features by linearizing only available features, skipping missing ones. In particular, if a feature $k$ is not available for the data sample $\mathbf{x}_{i,t}$, then the pair $c^{k}_{i,t}:\mathbf{x}^{k}_{i,t}]$ will be skipped and not included in the input to the LLM. This enables the use of a combination of heterogeneous data samples with different feature sets in both training and testing processes. This also allows the subsequent semantic recognition to proceed seamlessly on the generated text without the need for manual adjustments to account for data irregularities.